Find the mass in grams of Bi, atomic weight 208.9804 amu.
1
Expert's answer
2016-09-15T12:25:03-0400
The mass of the Bi can be calculated using the formula:
m(Bi) = M(Bi) × n(Bi)
Ar(Bi) = 208.9804 amu

M(Bi) = 208.9804 g/mol
The problem statement does not say the number of moles of the Bi, so we can suppose that this mean 1 mol. So the mass of the Bi will be:
m(Bi) = M(Bi) × n(Bi) = 208.9804 g/mo × 1 mol = 208.9804 g


Answer: 208.9804 g